# TAT-LLM: A Specialized Language Model for Discrete Reasoning over Tabular and Textual Data

### v1

@@ Abstract.
In this work, we address question answering (QA) over a hybrid of tabular and textual data that are very common content on the Web (e.g. SEC filings), where discrete reasoning capabilities are often required. Recently, large language models (LLMs) like GPT-4 have demonstrated strong multi-step reasoning capabilities. We then consider harnessing the amazing power of LLMs to solve our task. We abstract a Step-wise Pipeline for tabular and textual QA, which consists of three key steps, including Extractor, Reasoner and Executor, and initially design an instruction to instantiate the pipeline and validate that GPT-4 outperforms all existing methods. However, utilizing an online LLM like GPT-4 holds various challenges in terms of cost, latency, and data security risk, which motivates us to specialize smaller LLMs in this task. We develop a TAT-LLM language model by fine-tuning LLaMA 2 with the training data generated automatically from existing expert-annotated datasets following the Step-wise Pipeline. The experimental results have verified that our TAT-LLM model can outperform all baseline models, including the previous best fine-tuned models and very large-scale LLMs like GPT-4 on FinQA, TAT-QA and TAT-DQA benchmarks. We hope our work can serve as a pioneering example of specializing smaller language models for specific tasks.
Question Answering, Large Language Model, Tabular and Textual Data, Discrete Reasoning, Natural Language Processing
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Question answeringccs: Computing methodologies Natural language processing
1. Introduction
@@
Figure 1. Examples of QA with discrete reasoning over a hybrid of tabular and textual data.
@@
Nowadays the Web inundates its users with vast and intricate content, necessitating content analysis techniques. In this work, we focus on analyzing a prevalent category of Web content: documents containing tables and text, including but not limited to SEC filings, financial statements, academic papers, and medical reports. These documents often feature extensive numerical data in both tabular and textual formats, demanding discrete reasoning for intelligent comprehension. Recent research (Zhu et al., 2021; Chen et al., 2021; Zhu et al., 2022) investigates the intelligent comprehension of such documents through question answering (QA) tasks, as exemplified in Figure 1. The model, provided with a table and relevant text as the context, needs to perform various types of discrete reasoning, such as arithmetic calculations, making comparisons, and counting, to answer the question.
@@
To perform QA over hybrid tabular and textual data, a straightforward approach (Ran et al., 2019; Lei et al., 2022) involves taking the table, text, and question as input and generating the answer directly. However, this approach can be ineffective due to the complex reasoning process involved (Wei et al., 2022c). To address this, some works (Chen et al., 2021; Zhu et al., 2023a; Zhou et al., 2022; Li et al., 2023) decompose the task into multiple steps, producing intermediate results that serve as references for the final answer. For example, SoarGraph (Zhu et al., 2023a) first extracts the relevant information by selecting nodes in a constructed table-text graph structure and then employs a tree-based method to explicitly generate the mathematical expression of the answer. These multi-step approaches typically require distinct modules for each step and often optimize them concurrently through multi-task learning. However, the problem of how to decompose and model the answer process remains an open challenge.
@@
We identify three key steps in the process of tabular and textual QA, and abstract a Step-wise Pipeline, as illustrated in Figure 2 b). In particular, 1) Extractor identifies the relevant information or evidence to the question from the given context; 2) Reasoner generates a mathematical equation or logic rule with the obtained information; and 3) Executor derives the final answer by executing the mathematical equation or logic rule with the associated information. These three steps emphasize different abilities for tabular and textual QA — understanding of the question and context, knowledge (e.g., logic) of answering the question, and precision in calculating the answer. As these steps produce a sequence of intermediate results, we can specifically model them and enhance their discrete reasoning ability given a practical scenario of tabular and textual QA problems.
@@
A large language model (LLM) like GPT-3 (Brown et al., 2020) or FLAN (Wei et al., 2022a) has exhibited strong abilities required in the steps of the Step-wise Pipeline and an emergent multi-step reasoning ability (Wei et al., 2022b) with proper instructions such as chain-of-thought (CoT) (Wei et al., 2022c) and least-to-most (Zhou et al., 2023). We initially design a task instruction to instantiate the pipeline and validate that GPT-4 (OpenAI, 2023) outperforms existing methods for such QA tasks on multiple benchmarks (see Tables 24). However, utilizing an online LLM presents challenges in terms of cost, latency, and data security risks. On the other hand, fine-tuning a smaller language model specifically for math word problems has shown promising results (Cobbe et al., 2021a; Fu et al., 2023). This inspires us to explore the specialization of smaller language models for discrete reasoning over tabular and textual data within the Step-wise Pipeline.
@@
To this end, we develop a TAT-LLM model by fine-tuning LLaMA 2 (Touvron et al., 2023b) with the training data generated automatically from existing expert-annotated datasets, as shown in Figure 2 c). In particular, we train the selected LLM to take the table, text, and question as input, and complete the task in three steps following the Step-wise Pipeline. We construct each of the training instances with five parts: Instruction, Table, Text, Question, and Response, from available tabular and textual QA datasets. It is worth mentioning that the model output (or response) is formatted to a structured table, with each row corresponding to one step of the pipeline, as shown in the right part of Figure 2 c). This design is friendly to further refining the final answer and automatic evaluation. Moreover, to enhance the model performance, we equip the TAT-LLM model with an External Executor, which strengthens the execution of logical rules and mathematical calculations to better infer the final answer. We test our TAT-LLM on three popular tabular and textual QA benchmarks: FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021) and TAT-DQA (Zhu et al., 2022). The experimental results show that our smallest model TAT-LLM (7B) can outperform all baseline models and even beat GPT-4 on all three datasets.
We make the following main contributions in this work:
@@
  • We abstract a Step-wise Pipeline including Extractor, Reasoner and Executor for better discrete reasoning capabilities and ground the pipeline on LLMs, showcasing their potential on the tabular and textual QA task.
@@
  • We develop a TAT-GPT model by fine-tuning LLaMA 2 (Touvron et al., 2023b) for tabular and textual QA using the Step-wise Pipeline, which supports better for practical usage with cost and privacy concerns.
@@
  • We conduct extensive experiments on three benchmarks, which validate the superiority of our TAT-LLM model over both conventional methods and very large-scale LMs like GPT-4, and the scaling law of LLMs on the tabular and textual QA task.
@@
Figure 2. Comparison between a) End-to-end Pipeline and b) Step-wise Pipeline. c) Our TAT-LLM  language model is developed by fine-tuning LLaMA 2 (Touvron et al., 2023b) following the Step-wise Pipeline.
An example of QA with numerical reasoning over financial table-text data
Figure 2. Comparison between a) End-to-end Pipeline and b) Step-wise Pipeline. c) Our TAT-LLM  language model is developed by fine-tuning LLaMA 2 (Touvron et al., 2023b) following the Step-wise Pipeline.
2. Related Work
In this section, we review the related work on tabular and textual QA, large language model, and instruction tuning.
2.1. Tabular and Textual QA
@@
Documents containing a hybrid of tabular and textual contents are very common on the Web (Zhu et al., 2021). The intelligent comprehension of these documents is often achieved with QA tasks. Early works (Chen et al., 2020b; Li et al., 2021, 2021; Chen et al., 2020a) focus on answer extraction from tabular and textual data. For example, HybridQA (Chen et al., 2020b) requires the model to extract a cell or a text span as the answer to a given question from Wikipedia tables and text. In recent two years, a growing body of works tackle tabular and textual QA by performing discrete reasoning as massive numerical information is usually included in the table or text, such as TAT-QA (Zhu et al., 2021), TAT-HQA (Li et al., 2022a), FinQA (Chen et al., 2021), MultiHiertt (Zhao et al., 2022), and TAT-DQA (Zhu et al., 2022). In this line of research, many supervised fine-tuning methods are proposed, such as TagOp (Zhu et al., 2021), DyRRen (Li et al., 2023), ELASTIC (Zhang and Moshfeghi, 2022), MVGE (Wei et al., 2023), RegHNT (Lei et al., 2022), SoarGraph (Zhu et al., 2023a) and UniRPG (Zhou et al., 2022). For example, TagOp (Zhu et al., 2021) first adopts sequence tagging to extract the supporting evidence from the table and text, and then applies pre-defined operators like addition, subtraction, and counting to infer the final answer. Harnessing the capabilities of advanced LLMs to tackle tabular and textual QA remains a relatively underexplored area, which is the main focus of our work.
2.2. Large Language Models (LLMs)
@@
Recently, LLMs like ChatGPT (Ouyang et al., 2022a), GPT-4 (OpenAI, 2023) and PaLM (Chowdhery et al., 2022) have attracted tremendous research attention with their superb language understanding and generation abilities, bringing excellent performance to tasks of reading comprehension (Rogers et al., 2023), discrete reasoning(Wei et al., 2022c; Hendrycks et al., 2021; Cobbe et al., 2021b), etc. Since these general LLMs are usually extremely large and not accessible for most researchers, open-source smaller LLMs are proposed to allow researchers to customize their own LLMs for different tasks, such as LLaMA (Touvron et al., 2023a), Vicuna (Chiang et al., 2023), and Alpaca (Taori et al., 2023), with comparable performance to top general LLMs.
Existing research has attempted to instruct LLMs to perform table-based QA (Chen, 2023). Several works (Ye et al., 2023; Gemmell and Dalton, 2023; Jiang et al., 2023; Cheng et al., 2022) leverage general LLMs to solve the problem through a series of designed instructions or steps, which may include complex question decomposition, table evidence extraction, and SQL/code generation which is then executed via corresponding executors. However, these solutions do not focus on hybrid tabular and textual QA tasks, and rely on close-source general LLMs and carefully human-crafted instruction steps. Liu et al. (2023) focus on the multi-modal tabular, textual and image QA task, but still rely on instructing the general LLM to solve table-based questions. Our solution aims to enhance the tabular and textual QA ability by fine-tuning open-source LLMs.
2.3. Instruction Tuning
Instruction tuning is a very effective approach to boost the performance of pre-trained LLMs in various natural language tasks (Ouyang et al., 2022a; Zhang et al., 2023, 2023). It is an essential training stage for nearly all state-of-the-art LLMs (e.g., GPT-4 (OpenAI, 2023), Flan-T5 (Chung et al., 2022)), enabling pre-trained LLMs to understand the user query intent and follow user instructions. To perform instruction tuning, the LLM is trained on (input, output) pairs of natural language task data, where the input is usually a concatenation of task instruction, optional in-context examples and task input (Ouyang et al., 2022a). For general LLMs, instruction tuning usually aggregates massive data from a wide variety of natural language tasks (Chung et al., 2022) to develop well-round language abilities.
Apart from training general LLMs, some works also perform domain-specific instruction tuning to obtain LLMs for certain domains or tasks. For example, in financial domain, BloombergGPT (Wu et al., 2023) performs instruction tuning on a wide range of financial task data, largely outperforming general LLMs on financial tasks. Goat (Liu and Low, 2023) fine-tunes LLaMA on various arithmetic tasks, outperforming GPT-4 on some of these tasks. Model specialization (Fu et al., 2023) performs instruction tuning on small LMs to distill a certain task ability from general large LMs. Yang et al. (2023) distill table-based reasoning abilities from large LMs to smaller LMs. Similarly, our TAT-LLM focuses on financial tabular and textual QA tasks. To perform instruction tuning, techniques such as LoRA (Hu et al., 2022) and Qlora (Dettmers et al., 2023) are usually adopted to reduce the training cost.
3. Approach
@@
In this section, we first introduce the Step-wise Pipeline for tabular and textual QA to highlight its merits, and then provide details of our TAT-LLM model specializing in discrete reasoning over hybrid tabular and textual data.
3.1. End-to-end Pipeline vs. Step-wise Pipeline
We consider such a problem: given a question and its relevant context, the model aims to derive the answer from the context, where discrete reasoning capabilities are usually required. In this work, the context refers to the documents with a hybrid of tabular and textual data.
3.1.1. End-to-end Pipeline
As shown in Figure 2 a), to solve the problem, the model that follows an End-to-end Pipeline takes the given question and context as input and directly outputs the answer. This straightforward approach provides little insight into how the answer is derived. It is akin to a black-box method where the internal workings remain largely concealed.
3.1.2. Step-wise Pipeline
We abstract this pipeline from previous multi-step methods such as (Zhu et al., 2021, 2023a; Li et al., 2023). As shown in Figure 2 b), this pipeline addresses the task with three steps: Extractor, Reasoner, and Executor. In particular, Extractor serves as an information extraction module that identifies the relevant snippets or segments of information to the question from the context; 2) Reasoner works as a “logic thinker” to generate the right logic rule or equation that leads to the right answer to the question; 3) Executor is responsible for deriving the final answer by performing the logic rule or calculating the equation.
Such a Step-wise Pipeline is able to produce a sequence of intermediate results, which enables us to boost each component within the pipeline, consequently leads to improved final performance. For example, in this work, we choose to strengthen the third step by adding an extra Executor, which brings noticeable performance gains as verified experimentally in Section 4.6.3.
3.1.3. Implementing Pipelines with LLMs
LLMs are shown to be able to follow user instructions in natural language to generate desired results (Ouyang et al., 2022b). It therefore has become a general consensus that LLMs can be trained to complete specific tasks as expected with appropriate instructions. In this work, our goal is to instruct LLMs to perform discrete reasoning over tabular and textual data following the Step-wise Pipeline. To achieve it, we carefully design a natural language instruction to include the three steps in the Step-wise Pipeline, i.e., Extractor, Reasoner, and Executor. After that, we generate the training data with publicly available tabular and textual QA datasets and train LLMs to answer the questions to follow the instructions. At below, we elaborate on how we develop our TAT-LLM language model following the Step-wise Pipeline, and provide a comprehensive comparison with the model trained following the End-to-end Pipeline in Section 4.6.2.
3.2. TAT-LLM
Here we introduce how we develop the TAT-LLM model.
3.2.1. Selection of the Language Model
@@
We develop our TAT-LLM  by fine-tuning LLaMA 2  (Touvron et al., 2023b) on the training data transformed from publicly available expert-annotated tabular and textual QA datasets. LLaMA (Touvron et al., 2023a, b) is a series of open-source large language models that are trained on the large-scale publicly available corpus. It is commonly employed in the development of large language models with instruction tuning, such as Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), and Goat (Liu and Low, 2023). In this study, we choose the latest LLaMA 2 (Touvron et al., 2023b) as our base language model because it offers an extended maximum sequence length (i.e., $4,096$) compared to the previous version — LLaMA 1 (Touvron et al., 2023a), which often fails to accommodate the tabular and textual content of one instance from the existing QA datasets. Moreover, LLaMA 2 has demonstrated remarkable performance across various benchmarks.
@@
3.2.2. Construction of Training Data
@@
To develop TAT-LLM, our fine-tuning is based on the training data we automatically transform from available tabular and textual QA datasets. For implementation in this work, we use three datasets, including FinQA (Chen et al., 2021), TAT-QA(Zhu et al., 2021) and TAT-DQA(Zhu et al., 2022), which are detailed in Section 4.1. We carefully design templates for each dataset to generate the training data for fine-tuning LLMs. Please refer to Appendix A for detailed information about their formats. As shown in Figure 2 c), each training instance is designed to be composed of five parts, including Instruction, Table, Text, Question, and Response.
@@
Instruction: This part provides a detailed guide for what the task is and how to complete it. It outlines the steps that should be followed to derive the answer from the provided context. We adopt the Step-wise Pipeline described in above sections. 1) Extractor: We request the model to extract the relevant information from either the table or the accompanying text. 2) Reasoner: Based on the extracted information, we ask the model to derive an equation or a logic rule that is used to infer the answer. 3) Executor: We require the model to execute the equation or logic rule to arrive at the final answer. An illustration is given in Figure 2 c), which is the same for all the training instances. The instruction part helps the model to precisely understand the task and generate accurate and meaningful outputs.
@@
Table: This part refers to the tabular data in the input context of this task. We compose this part using the tables from FinQA and TAT-QA datasets. As each table in the two datasets is stored in a Two-Dimensional (2D) array, we transform it into a markdown table. Note that we omit TAT-DQA dataset for constructing this part, considering the context given in TAT-DQA is the document pages from PDF documents, where the position and structure of the table are unknown.
@@
Text: This part refers to the textual data in the input context. For FinQA and TAT-QA datasets, we compose this part with the annotated paragraphs associated with the table; all contents from the document pages in TAT-DQA dataset are regarded as textual data.
@@
Question: This part refers to the question that is asked based on the input tabular and textual data. We directly put the question from the three datasets in this part. The goal of the task is to generate the right answer to this question.
Response: This part describes the model’s output based on the given table, text, and question. To better inspect the results in the intermediate steps and automatically evaluate the model output, we train the language model to format the output in a structured markdown table with two columns, i.e., step and output. Each row in the output table corresponds to one step defined in the instruction. In the end, the model draws its conclusion with a statement: ”The answer is: answer”, making its final response clear.
To train the model, the response part in the training set needs to be generated in advance. Existing tabular and textual QA datasets often provide the answer to the question and the derivation annotated by human experts. To generate the correct response for constructing our training instance, we first automatically identify the supporting evidence from the derivation of each question, such as the numbers and entities, which are used as the output of the first step. We then convert the derivation to a valid equation or logic rule as the output of the second step, making sure its execution result is consistent with the answer. We directly use the annotated answer as the output of the last step.
3.2.3. Training
After constructing the training set, we train our TAT-LLM model in various sizes, including 7B, 13B, and 70B, by fine-tuning LLaMA 2 (Touvron et al., 2023b) using Low-Rank Adaptation (LoRA) (Hu et al., 2022). More details are provided in Section 4.3.
3.2.4. External Executor
We observe that the trained model struggles in performing the execution of the mathematical equations or logic rules (e.g., counting, comparison) in the last step, according to experimental evaluations in Section 4.6.3. Such incompetence of the Executor has also been found in previous works (Luo et al., 2023; Yue et al., 2023; Liu and Low, 2023). To enhance the model’s accuracy of the final output, we propose to add an External Executor to our model, which takes the model’s intermediate outputs as input and performs the execution of the equations or logic rules to obtain the final answer. With this External Executor, the model is able to refine its output to a better one. With the output of the model formatted to a structured table, as aforementioned, we can easily access the intermediate results. The whole process is summarized in Algorithm 1. After obtaining the intermediate results, the External Executor is applied to refine the final answer instead of directly using the prediction of the model.
4. Experiments
We conduct extensive experiments to evaluate the effectiveness of our Step-wise Pipeline and TAT-LLM model. In addition to the comparisons with previous work, we also provide a comprehensive analysis to further verify our model design rationality, and reveal more properties of our method.
4.1. Datasets
We conduct the experiments using three publicly available tabular and textual QA datasets, i.e., FinQA (Chen et al., 2021), TAT-QA (Zhu et al., 2021) and TAT-DQA (Zhu et al., 2022), which are introduced at below. See Table 1 for the statistics of splits of each dataset.
FinQA (Chen et al., 2021) is an expert-annotated tabular and textual QA dataset in which the tables and text are sampled from financial reports. It focuses on discrete reasoning capabilities like addition, subtraction, multiplication, division, and numerical comparison.
TAT-QA (Zhu et al., 2021) is also built with tables and paragraphs extracted from financial reports. Most questions require discrete reasoning for generating the answers, and meanwhile there are also cases where the answers can be extracted directly from the tables or text. Its questions are classified into four different types: Span, Multiple Spans, Counting, and Arithmetic. In addition to the numerical value in the predicted answer, it also requires the model to predict the right scale of the answer, e.g., thousand, million.
TAT-DQA (Zhu et al., 2022) is an extension of the TAT-QA dataset, focusing on question answering over the original long financial statements with up to three pages, and the position and structure of the tables are unknown. This setting is believed to be more realistic and challenging than FinQA and TAT-QA datasets.
Table 1. Statistics of the train, validation and test splits for the financial tabular and textual QA datasets.

### v3

@@ Abstract
In this work, we address question answering (QA) over a hybrid of tabular and textual data, involving a variety of common content in reality like SEC filings, where discrete reasoning is often required. We consider harnessing the multi-step reasoning capabilities of large language models (LLMs) to tackle this problem, which have recently achieved remarkable success in many natural language tasks. To do this, we first abstract a Step-wise Pipeline for tabular and textual QA to help LLMs better execute multi-step inference, containing three key steps of Extractor, Reasoner and Executor. We initially design an instruction to validate the pipeline on GPT-4, demonstrating promising results. However, utilizing an online LLM like GPT-4 holds various challenges in terms of cost, latency, and data security risk, which motivates us to specialize smaller LLMs in this task. We then develop a TAT-LLM model by fine-tuning LLaMA 2 with the training data generated automatically from existing datasets following the Step-wise Pipeline. The experimental results have verified that our TAT-LLM model can outperform all compared models, including prior best fine-tuned models and very large-scale LLMs like GPT-4 on FinQA, TAT-QA and TAT-DQA benchmarks. The models and datasets are publicly available at https://huggingface.co/next-tat
1 Introduction
@@
Figure 1: Examples of QA with discrete reasoning over a hybrid of tabular and textual data.
@@
The documents containing both tables and text, e.g. SEC filings, academic papers and medical reports, make a very prevalent category of content in the real world. They often feature extensive numerical data in both the tabular and textual content, necessitating discrete reasoning capabilities for machines to comprehend them. Recent research Zhu et al. (2021); Chen et al. (2021) investigates the intelligent comprehension of such documents through question answering (QA) tasks, as exemplified in Figure 1. The model, provided with a table and relevant text as the context, needs to perform various types of discrete reasoning, such as arithmetic calculations, making comparisons, and counting, to answer the question.
@@
To perform QA over hybrid tabular and textual data, a straightforward approach Ran et al. (2019) involves taking the table, text, and question as input and generating the answer directly. This approach can be ineffective due to the complex reasoning process involved Wei et al. (2022c). To address this issue, some works Lei et al. (2022); Zhou et al. (2022); Zhu et al. (2023) decompose the task into multiple steps, producing intermediate results that serve as references for the final answer. These multi-step approaches typically design distinct modules at each step and often optimize these modules concurrently through multi-task learning. To date, there has been no consensus on how to decompose the answer process in existing literature.
@@
Recently, large language models (LLMs) like GPT-4 OpenAI (2023) and FLAN Wei et al. (2022a) have exhibited strong multi-step reasoning abilities Wei et al. (2022b) with proper instructions such as chain-of-thought (CoT) Wei et al. (2022c) and least-to-most Zhou et al. (2023). Therefore, we consider harnessing this amazing power of LLMs for better discrete reasoning over hybrid tabular and textual data. To achieve this, we first identify three key steps in the process of tabular and textual QA from previous multi-step methods Zhu et al. (2021, 2023); Zhou et al. (2022), and abstract a Step-wise Pipeline, as illustrated in Figure 2 b). Specifically, 1) Extractor identifies the relevant information or evidence to the question from the given context; 2) Reasoner generates a mathematical equation or logic rule with the obtained information; and 3) Executor derives the final answer by executing the mathematical equation or logic rule with the associated information. The three steps emphasize different capabilities of the tabular and textual QA model — understanding the question and context, inferring the logic for answering the question, and calculating the answer with precision. These steps produce a sequence of intermediate results, which means we can specifically model and enhance one (or more) of them given a specific application scenario.
@@
Following the Step-wise Pipeline, we initially design a task instruction and validate it on GPT-4 OpenAI (2023), achieving promising results on multiple benchmarks. However, utilizing an online LLM presents challenges in terms of cost, latency, and data security risk. By contrast, fine-tuning a smaller language model, specifically for math word problems Fu et al. (2023); Cobbe et al. (2021), has been proven fairly appealing. We are then motivated to explore the specialization of smaller language models for addressing this challenge following the Step-wise Pipeline.
@@
We develop a TAT-LLM model by fine-tuning LLaMA 2 Touvron et al. (2023b) with the training data generated automatically from existing expert-annotated datasets, as shown in Figure 2 c). In particular, we train the selected LLM to take the table, text, and question as input, and complete the task in three key steps following the Step-wise Pipeline. We construct each of the training instances with five parts: Instruction, Table, Text, Question, and Response, from available tabular and textual QA datasets. Inspired by Tab-CoT Ziqi and Lu (2023), the model output (or response) is formatted to a structured table, with each row corresponding to one step of the pipeline, as shown in the right part of Figure 2 c). This design is friendly to further refining the outputs at each step and automatic evaluation. Moreover, to enhance the model performance, we equip the TAT-LLM model with an External Executor, which strengthens the execution of logical rules and mathematical calculations to better infer the final answer. We test our TAT-LLM on three popular benchmarks: FinQA Chen et al. (2021), TAT-QA Zhu et al. (2021) and TAT-DQA Zhu et al. (2022). The experimental results show that our smallest model TAT-LLM (7B) can outperform all baseline models and even beat GPT-4 on all three datasets. In summary, we make the following main contributions in this work:
@@
  • We abstract a Step-wise Pipeline including Extractor, Reasoner and Executor to assist LLMs in better performing discrete reasoning over a hybrid of tabular and textual data.
@@
  • We develop a TAT-LLM model by fine-tuning LLaMA 2 (including $7$B, $13$B and $70$B) for tabular and textual QA following the Step-wise Pipeline, better supporting practical application with cost and privacy concerns.
@@
  • We conduct extensive experiments on three popular benchmarks, validating the superiority of our TAT-LLM model over both conventional methods and very large-scale LMs, e.g. GPT-4.
@@
Figure 2: Comparison between a) End-to-end Pipeline and b) Step-wise Pipeline. c) Our TAT-LLM  language model is developed by fine-tuning LLaMA 2 following the Step-wise Pipeline.
2 Related Work
2.1 Tabular and Textual QA
@@
Early works on tabular and textual QA Chen et al. (2020b); Li et al. (2021); Chen et al. (2020a) focus on answer extraction from tabular and textual data. In recent two years, a growing body of works tackle tabular and textual QA by performing discrete reasoning as massive numerical information is usually included in the table or text, such as TAT-QA Zhu et al. (2021), TAT-HQA Li et al. (2022), FinQA Chen et al. (2021), MultiHiertt Zhao et al. (2022), and TAT-DQA Zhu et al. (2022). In this line of research, many supervised fine-tuning methods are proposed, such as DyRRen Li et al. (2023), RegHNT Lei et al. (2022), UniRPG Zhou et al. (2022) and MVGE Wei et al. (2023). Harnessing the capabilities of advanced LLMs to tackle tabular and textual QA remains a relatively underexplored area, which is the main focus of our work.
2.2 Large Language Models (LLMs)
@@
Recently, LLMs like ChatGPT Ouyang et al. (2022) have attracted tremendous research attention with their superb language understanding and generation abilities, bringing excellent performance to many tasks like reading comprehension Rogers et al. (2023), discrete reasoning Wei et al. (2022c), etc. Since these general LLMs are usually extremely large and not accessible for most researchers, open-source smaller LLMs are proposed to allow researchers to customize their own LLMs for different tasks, such as LLaMA Touvron et al. (2023a) and Alpaca Taori et al. (2023), with comparable performance to the general LLMs. Recently, there has been a trend in the research community that small LLMs are specialized for specific tasks through fine-tuning with instructions Fu et al. (2023), yielding impressive performance. To the best of our knowledge, we are the first work to specialize a small LLM (e.g.,7B) in tabular and textual QA challenge.
3 Approach
@@
In this section, we first introduce the Step-wise Pipeline for tabular and textual QA, and then elaborate our proposed TAT-LLM model.
3.1 Step-wise Pipeline
Large language models (LLMs) like ChatGPT Ouyang et al. (2022) and FLAN Wei et al. (2022a) have demonstrated astoundingly strong multi-step reasoning abilities Wei et al. (2022b) following human instructions in many tasks Zhou et al. (2023); Cobbe et al. (2021). To better leverage such amazing power of LLMs to address QA over hybrid tabular and textual data, we abstract a Step-wise Pipeline from the previous multi-step methods Zhu et al. (2021); Zhou et al. (2022) on this task. As shown in Figure 2 b), this pipeline addresses the task with three key steps: Extractor, Reasoner, and Executor. In particular, Extractor serves as an information extraction module that identifies the relevant snippets or segments of information to the question from the context; 2) Reasoner works as a “logic thinker” to generate the right logic rule or equation that leads to the right answer to the question; 3) Executor is responsible for deriving the final answer by performing the logic rule or executing the equation. Compared with an End-to-end Pipeline that provides little insight into how the answer is derived, as shown in Figure 2 a), such a Step-wise Pipeline is able to produce a sequence of intermediate results, which means we can boost each component (or some of them) within the pipeline to consequently acquire improved final performance. For example, in this work, we choose to strengthen the third step by adding an extra Executor, which brings noticeable performance gains as verified experimentally in Section 4.3.
To instruct LLMs to perform discrete reasoning over tabular and textual data following the Step-wise Pipeline, we carefully design a natural language instruction to include the three key steps in the Step-wise Pipeline. Please refer to Appendix D for more information. We validate it on both FinQA and TAT-QA benchmarks using ChatGPT Ouyang et al. (2022) and GPT-4 OpenAI (2023) and obtain promising results, demonstrating the rationality of applying the Step-wise Pipeline grounded on the LLMs. We then generate training data with publicly available tabular and textual QA datasets and specialize a smaller LLM, i.e., LLaMA 2 Touvron et al. (2023b), to answer questions following the Step-wise Pipeline for better supporting practical applications with cost and privacy concerns, as elaborated at below.
3.2 TAT-LLM
Selection of Language Model.
@@ 3.2 TAT-LLM
We develop our TAT-LLM  by fine-tuning LLaMA 2  Touvron et al. (2023b). LLaMA Touvron et al. (2023a) is a series of open-source large language models trained on the large-scale publicly available corpus. It is commonly employed in the development of large language models with instruction tuning, such as Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023). In this study, we choose the latest LLaMA 2 Touvron et al. (2023b) as our base language model considering it offers an extended maximum sequence length (i.e., $4,096$) compared to the previous version LLaMA Touvron et al. (2023a), which often fails to accommodate the tabular and textual content of one instance from existing datasets. Moreover, LLaMA 2 has demonstrated remarkable performance across various benchmarks.
@@
Construction of Training Data.
@@
Our fine-tuning is based on the training data we automatically transform from existing tabular and textual QA datasets. For implementation, we use three datasets, i.e., FinQA Chen et al. (2021), TAT-QA Zhu et al. (2021) and TAT-DQA Zhu et al. (2022). We carefully design templates for each dataset to generate the training data for fine-tuning LLMs. Please refer to Appendix C for more information. As shown in Figure 2 c), each training instance is designed to be composed of five parts, including Instruction, Table, Text, Question, and Response.
@@
$\bullet$ Instruction: This part provides a detailed guide for what the task is and how to complete it. It outlines the steps that should be followed to derive the answer from the provided context. We adopt the Step-wise Pipeline described in above sections. 1) Extractor: We request the model to extract the relevant information from either the table or the accompanying text. 2) Reasoner: Based on the extracted information, we ask the model to derive an equation or a logic rule that is used to infer the answer. 3) Executor: We require the model to execute the equation or logic rule to arrive at the final answer. An illustration is given in Figure 2 c), which is the same for all the training instances. The instruction part helps the model to precisely understand the task and generate accurate and meaningful outputs.
@@
$\bullet$ Table: This part refers to the tabular data in the input context of this task. We compose this part using the tables from FinQA and TAT-QA datasets. As each table in the two datasets is stored in a Two-Dimensional (2D) array, we transform it into a markdown table. Note that we omit TAT-DQA dataset for this part, considering the context given in TAT-DQA is the document pages from PDF files, where the structure of the table is unknown.
@@
$\bullet$ Text: This part refers to the textual data in the input context. For FinQA and TAT-QA datasets, we compose this part with the paragraphs relevant to the table; all contents from the document pages in TAT-DQA dataset are regarded as textual data.
@@
$\bullet$ Question: This part refers to the question that is asked based on the given tabular and textual context. We directly put the question from the three datasets in this part. The goal of the task is to generate the right answer to this question.